Write an equation in point-slope form of the line that passes through the given point and has the given slope m:
(5,-12), m = -2/5
(thats two over five)
Would you please also work out the problem so i know and understand how to do it in the future?

Answer by TimothyLamb(4379)

---
point-slope form:
y - y1 = m(x - x1)
---
answer:
y + 12 = (-2/5)(x - 5)
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php